J. Äystö,^{1} A. Badertscher,^{2} L. Bezrukov,^{3} J. Bouchez,^{4} A. Bueno,^{5} J. Busto,^{6} J.-E. Campagne,^{7} Ch. Cavata,^{8} A. de Bellefon,^{9} J. Dumarchez,^{10} J. Ebert,^{11} T. Enqvist,^{12} A. Ereditato,^{13} F. von Feilizsch,^{14} M. Göger-Neff,^{14} S. Gninenko,^{15} W. Gruber,^{2} C. Hagner,^{11} K. Hochmuth,^{16} J. Holeczek,^{17} J. Kisiel,^{17} L. Knecht,^{2} I. Kreslo,^{13} V. A. Kudryavtsev,^{18} P. Kuusiniemi,^{12} T. Lachenmaier,^{15} M. Laffranchi,^{2} B. Lefievre,^{9} M. Lindner,^{19} J. Maalampi,^{20} A. Marchionni,^{2} T. Marrodn Undagoitia,^{14} A. Meregaglia,^{2} M. Messina,^{13} M. Mezzetto,^{21} L. Mosca,^{15} U. Moser,^{13} A. Mller,^{2} G. Natterer,^{2} L. Oberauer,^{14} P. Otiougova,^{2} T. Patzak,^{9} J. Peltoniemi,^{22} W. Potzel,^{14} C. Pistillo,^{13} G. G. Raffelt,^{16} M. Roos,^{23} B. Rossi,^{13} A. Rubbia,^{2} N. Savvinov,^{13} N. Spooner,^{18} A. Tonazzo,^{9} W. Trzaska,^{1} J. Ulbricht,^{2} C. Volpe,^{24} J. Winter,^{14} M. Wurm,^{14} A. Zalewska,^{15} and R. Zimmermann^{11}

1Department of Physics, University of Jyväskylä, Finland 2Institut für Teilchenphysik, ETHZ, CH-8093 Zürich, Switzerland 3Institute for Nuclear Research, Russian Academy of Sciences, Moscow, 117312 Russia 4DAPNIA/SPP CEA-Saclay 91191 Gif-sur-Yvette cedex and APC, Paris 5Dpto Fisica Teorica y del Cosmos & C.A.F.P.E., Universidad de Granada, Spain 6Centre de Physique des Particules de Marseille (CPPM), IN2P3-CBRS et Universit d’Aix-Marseille II, 163 Av. de Luminy, B.P. 902, 13288 Marseille Cedex 09, France 7Laboratoire de l’Acclrateur Linaire (LAL), IN2P3-CNRS et Universit PARIS-SUD 11, Centre Scientifique d’Orsay, B.P. 34, 91898 ORSAY Cedex, France 8CEA - Saclay, 91191 Gif sur Yvette Cedex, France 9Astroparticule et Cosmologie (APC), UMR 7164 (CNRS, Universiti Paris VII, CEA, Observatoire de Paris), 11 pl. Marcelin Berthelot, 75231 Paris Cedex 05, France 10Laboratoire de Physique Nuclaire et des Hautes Energies (LPNHE), IN2P3-CNRS et Universits Paris VI et Paris VII, 4 place Jussieu, Tr. 33 - RdC, 75252 Paris Cedex 05, France 11Universitt Hamburg, Institut fr Experimentalphysik, Geb. 67/216, Luruper Chaussee 149, 22761 Hamburg, Germany 12CUPP, University of Oulu, Finland 13Laboratory for High Energy Physics, University of Bern, Sidlerstrasse, 5, CH-3012 Bern, Switzerland 14Technical University Munich, Physikdepartment E15, James-Franck-Str 1, 85748 Garching, Germany 15Unknown 16Max-Planck-Institut für Physik (Werner-Heisenberg-Institut), Föhringer Ring 6, 80805 München, Germany 17Institute of Physics, University of Silesia, Uniwersytecka 4, PL-40007 Katowice, Poland 18Particle Physics and Particle Astrophysics, Department of Physics and Astronomy, University of Sheffield, Hicks Building, Hounsfield Road, Sheffield S3 7RH, UK 19 Max-Planck-Institut fuer Kernphysik, Saupfercheckweg 1, D-69117 Heidelberg, Germany 20 University of Jyvskyl, Finland 21INFN-Univ. di Padova, Dept. di Fisica, via Marzolo 8, 35100 Padua, Italy 22 CUPP, University of Oulu, Finland 23Department of Physical Sciences, University of Helsinki, Finland 24Institut de Physique Nucleaire d’Orsay (IPNO), Groupe de Physique Theorique, Universite de Paris-Sud XI, Bat 100, rue Georges Clemenceau, 91406 Orsay, Cedex, France

(Dated: November 6, 2006)